OpenStudy (anonymous):

a vial contains 2000mg. which is to be used for 2 dosages. one patient is to be administered 660 more than another. how much should be administered to each?

OpenStudy (anonymous):

i got 670 and 1330 is this correct?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

x + x + 660 = 2000 2x + 660 = 2000 x + 330 = 1000 x = 1000 - 330

OpenStudy (anonymous):

670

OpenStudy (amistre64):

670 330 ------ 1000 i like it

OpenStudy (anonymous):

good

OpenStudy (amistre64):

670 660 ---- 1330

OpenStudy (anonymous):

yes that's what i get

OpenStudy (amistre64):

:)